# AnomalyDiffusion: Few-Shot Anomaly Image Generation with Diffusion Model

### 10 Dec 2023 v1

@@ 1 Introduction
Figure 1: Top: Our model generates extensive anomaly data, which supports the downstream Anomaly Detection (AD), Localization (AL) and Classification (AC) tasks, while previous methods mainly rely on unsupervised learning or few-shot supervised learning due to the limited anomaly data; Bottom: Generated anomaly results on hazelnut-crack and capsule-squeeze of our model and existing anomaly generation methods. Our model generates the most authentic anomalies over all the methods.
@@ 2.1 Generative Models
Few-shot image generation. Few-shot image generation aims to generate diverse images with limited training data. Early methods propose modifying network weights (Mo, Cho, and Shin 2020), using various regularization techniques (Li et al. 2020) and data augmentation (Tran et al. 2021) to prevent overfitting. To deal with the extremely limited data (less than 10), recent works  (Ojha et al. 2021; Hu et al. 2023) introduce new cross-domain consistency losses to keep the generated distribution. Textual Inversion  (Gal et al. 2022) and Dreambooth (Ruiz et al. 2023) encode a few images into the textual space of a pre-trained LDM, but are unable to control the generated locations accurately.
@@ 2.2 Anomaly Inspection
Anomaly generation. The scarcity of anomaly data has sparked research interest in anomaly generation. DRAEM (Zavrtanik, Kristan, and Skočaj 2021), Cut-Paste (Li et al. 2021), Crop-Paste (Lin et al. 2021) and PRN (Zhang et al. 2023a) crop and paste unrelated textures or existing anomalies into normal sample. But they either generate less realistic anomalies or have limited generated diversity. The GAN-based model SDGAN (Niu et al. 2020) and Defect-GAN (Zhang et al. 2021), generate anomalies on normal samples by learning from anomaly data. But they require a large amount of anomaly data and cannot generate anomaly mask. DMFGAN (Duan et al. 2023) transfers a StyleGAN2 (Karras et al. 2020) pretained on normal samples to anomaly domain, but lacks generation authenticity and accurate alignment between generated anomalies and masks. In contrast, our model incorporates spatial anomaly embedding and adaptive attention re-weighting mechanism, which can generate anomalous image-mask pairs with great diversity and authenticity.
@@ 4.1 Experiment Settings
Dataset. we conduct experiments on the widely used MVTec (Bergmann et al. 2019) dataset. We employ one-third of the anomaly data as the training set, reserving the remaining two-thirds for testing.
@@ 4.4 Ablation Study
We evaluate the effectiveness of our components: spatial anomaly embedding (SAE), masked diffusion loss (Masked $\mathcal{L}$) and adaptive attention re-weighting mechanism (AAR). Not that the models without SAE employs only an anomaly embedding trained by textual inversion. We train 5 models: 1) with none of these components; 2) only SAE; 3) SAE + masked $\mathcal{L}$; 4) masked $\mathcal{L}$ + AAR and 5) the full model (ours). We employ these models to generate 1000 anomalous image-mask pairs and train an U-Net for anomaly localization. We compare the pixel-level localization results in Tab. 6. It demonstrates that the omission of any of the proposed modules leads to a noticeable decline in the model’s performance on anomaly localization, which validates the efficacy of the proposed modules.
@@ 5 Conclusion
In this paper, we propose Anomalydiffusion, a novel anomaly generation model which generates anomalous image-mask pairs. We disentangle anomaly information into anomaly appearance and location information that are represented by anomaly embedding and spatial embedding in the textual space of latent diffusion separately. Moreover, we also introduce an adaptive attention re-weighting mechanism, which helps our model focus more on the areas with less noticeable generated anomalies, thus improving the alignment between the generated anomalies masks. Extensive experiments show that our model outperforms the existing anomaly generation methods and our generated anomaly data effectively improves the performance of the downstream anomaly inspection tasks. In future work, we would explore the application of a more potent diffusion model to enhance the resolution of the generated anomalies, where higher resolutions could potentially further imporve the results of our methodology.
Acknowledgement
This work was supported by National Natural Science Foundation of China (62302297, 72192821, 62272447), Young Elite Scientists Sponsorship Program by CAST (2022QNRC001), Shanghai Sailing Program (22YF1420300), Beijing Natural Science Foundation (L222117), the Fundamental Research Funds for the Central Universities (YG2023QNB17), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Science and Technology Commission (21511101200).
@@ Appendix A Overview
  • Ablation study on Spatial Anomaly Embedding (Sec. D).
@@ C.1 Training Details
Training spatial anomaly embedding. For each anomaly type, an anomaly embedding $e_{a}$ is assigned, while a shared spatial encoder $E$ is employed across all anomaly categories. Each anomaly embedding $e_{a}$ is composed of 8 tokens, and the spatial embedding $e_{s}$ comprises 4 tokens. The batch size is set at 4, and the learning rate is 0.005. During each training iteration, we randomly sample 4 anomalous image-mask pairs from all the anomaly categories. We train all the anomaly embedding and spatial encoder at the same time for 300K iterations.
@@
Appendix D Ablation on Spatial Anomaly Embedding
@@
We aim to seek a text embedding that guides the latent diffusion model in generating anomalies within a given anomaly mask. However, textual inversion tends to capture the location of anomalies along with the anomaly type information, which results in the generated anomalies only distributed in specific locations. Therefore, we propose spatial anomaly embedding $e$, consisting of an anomaly embedding $e_{a}$ (for appearance) and a spatial embedding $e_{s}$ (for location), which disentangles the spatial information from anomaly appearance. To further validate this theory, we directly employ text inversion to train an anomaly embedding and use the trained embedding to generate anomalies with a given mask by blended latent diffusion (the same as our generation process). The generated results are shown in Fig. 6. It can be seen that the generated result by Textual Inversion tends to generate anomalies at the location the same as the training sample, which limits its application in anomaly generation where anomalies can locate at arbitrary positions.
@@ Appendix G More Quantitative experiments
In this section, we provide supplementary experiments to complement those presented in the main paper. Specifically, in addition to the methods covered in the main paper, we include Crop&Paste (Lin et al. 2021) for comparison and we additionally introduce the Per Region Overlap (PRO) metric to provide a more comprehensive evaluation on anomaly localization. The experiment settings are the same as that in the main paper, where we train an U-net on the generated anomaly data. The pixel-level anomaly localization results are shown in Tab. 8 and the image-level anomaly detection results are shown in Tab. 9. The quantitative results demonstrate that our model outperforms all the other anomaly generation methods in terms of both anomaly localization and detection, indicating our good anomaly generation quality and diversity.

### 22 Feb 2024 v2

@@ 1 Introduction
Figure 1: Top: Our model generates extensive anomaly data, which supports the downstream Anomaly Detection (AD), Localization (AL) and Classification (AC) tasks, while previous methods mainly rely on unsupervised learning or few-shot supervised learning due to the limited anomaly data; Bottom: Generated anomaly results on hazelnut-crack and capsule-squeeze of our model and existing anomaly generation methods, where our results are the most authentic.
@@ 2.1 Generative Models
Few-shot image generation. Few-shot image generation aims to generate diverse images with limited training data. Early methods propose modifying network weights (Mo, Cho, and Shin 2020), using various regularization techniques (Li et al. 2020) and data augmentation (Tran et al. 2021) to prevent overfitting. To deal with the extremely limited data (less than 10), recent works  (Ojha et al. 2021; Wang et al. 2022; Hu et al. 2023) introduce cross-domain consistency losses to keep the generated distribution. Textual Inversion  (Gal et al. 2022) and Dreambooth (Ruiz et al. 2023) encode a few images into the textual space of a pre-trained LDM, but cannot control the generated locations accurately.
@@ 2.2 Anomaly Inspection
Anomaly generation. The scarcity of anomaly data has sparked research interest in anomaly generation. DRAEM (Zavrtanik, Kristan, and Skočaj 2021), Cut-Paste (Li et al. 2021), Crop-Paste (Lin et al. 2021) and PRN (Zhang et al. 2023a) crop and paste unrelated textures or existing anomalies into normal sample. But they either generate less realistic anomalies or have limited generated diversity. The GAN-based model SDGAN (Niu et al. 2020) and Defect-GAN (Zhang et al. 2021), generate anomalies on normal samples by learning from anomaly data. But they require a large amount of anomaly data and cannot generate anomaly mask. DFMGAN (Duan et al. 2023) transfers a StyleGAN2 (Karras et al. 2020) pretained on normal samples to anomaly domain, but lacks generation authenticity and accurate alignment between generated anomalies and masks. In contrast, our model incorporates spatial anomaly embedding and adaptive attention re-weighting mechanism, which can generate anomalous image-mask pairs with great diversity and authenticity.
@@ 4.1 Experiment Settings
Dataset. we conduct experiments on the widely used MVTec (Bergmann et al. 2019) dataset. We employ one-third of the anomaly data with the lowest ID numbers as the training set, reserving the remaining two-thirds for testing.
@@ 4.4 Ablation Study
We evaluate the effectiveness of our components: spatial anomaly embedding (SAE), masked diffusion loss (Masked $\mathcal{L}$) and adaptive attention re-weighting mechanism (AAR). Not that the models without SAE employ only an anomaly embedding trained by textual inversion. We train 5 models: 1) with none of these components; 2) only SAE; 3) SAE + masked $\mathcal{L}$; 4) masked $\mathcal{L}$ + AAR and 5) the full model (ours). We employ these models to generate 1000 anomalous image-mask pairs and train an U-Net for anomaly localization. We compare the pixel-level localization results in Tab. 6. It demonstrates that the omission of any of the proposed modules leads to a noticeable decline in the model’s performance on anomaly localization, which validates the efficacy of the proposed modules. For more experiments, please refer to the supplementary material.
@@ 5 Conclusion
In this paper, we propose Anomalydiffusion, a novel anomaly generation model which generates anomalous image-mask pairs. We disentangle anomaly information into anomaly appearance and location information represented by anomaly embedding and spatial embedding in the textual space of LDM. Moreover, we also introduce an adaptive attention re-weighting mechanism, which helps our model focus more on the areas with less noticeable generated anomalies, thus improving the alignment between the generated anomalies and masks. Extensive experiments show that our model outperforms the existing anomaly generation methods and our generated anomaly data effectively improves the performance of the downstream anomaly inspection tasks. In future work, we would explore the application of a more potent diffusion model to enhance the resolution of the generated anomalies, which could further improve the performance.
Acknowledgments
This work was supported by National Natural Science Foundation of China (62302297, 72192821, 62272447), Young Elite Scientists Sponsorship Program by CAST (2022QNRC001), Shanghai Sailing Program (22YF1420300), Beijing Natural Science Foundation (L222117), the Fundamental Research Funds for the Central Universities (YG2023QNB17), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Science and Technology Commission (21511101200), CCF-Tencent Open Research Fund (RAGR20220121).
@@ Appendix A Overview
  • More ablation studies (Sec. D).
@@ C.1 Training Details
Training spatial anomaly embedding. For each anomaly type, an anomaly embedding $e_{a}$ is assigned, while a shared spatial encoder $E$ is employed across all anomaly categories. Each anomaly embedding $e_{a}$ is composed of 8 tokens, and the spatial embedding $e_{s}$ comprises 4 tokens. The batch size is set at 4, and the learning rate is 0.005. During each training iteration, we randomly sample 4 anomalous image-mask pairs from all the anomaly categories. We train all the anomaly embedding and spatial encoder at the same time for 300K iterations in 3 days on an NVIDIA GeForce RTX 3090 24GB GPU.
@@
C.2 Metrics
In the quantitative experiments, we employ the following metrics to measure the model performance.
  • Inception Score(IS) quantifies the quality and diversity of generated images by computing the exponential of the negative of the KL divergence between the marginal distribution of generated images and the conditional distribution of class labels predicted by an Inception model. A higher IS score represents a better generation quality and diversity.
  • Intra-cluster Pairwise LPIPS Distance (IC-LPIPIS) (Ojha et al. 2021) clusters the generated images into $k$ groups based on LPIPS distance to $k$ target samples, and then compute the average mean LPIPS distances to corresponding target samples within each cluster. A higher IC-LPIPS indicates a better generation diversity.
  • Area Under the Receiver Operating Characteristic (AUROC) measures the performance of a binary classification model by evaluating its ability to distinguish between true positive and false positive rates across different probability thresholds. A higher AUROC means better anomaly detection and localization performance.
  • Average Precision (AP, which is also known as PR-AUC) assesses the precision-recall curve for a classification model, calculating the average precision of the model across different recall levels, providing a summary of its overall performance. A higher AP means better anomaly detection and localization results.
  • $\mathbf{F_{1}}$-max is a variant of the $F_{1}$ score that maximizes both precision and recall by selecting the threshold that yields the highest F1 score when evaluating a binary classification model. A higher $\mathbf{F_{1}}$-max represents better anomaly detection and localization results
Appendix D More Ablation Studies
D.1 Ablation on Spatial Anomaly Embedding
@@
We aim to seek a text embedding that guides the latent diffusion model in generating anomalies within a given anomaly mask. However, textual inversion tends to capture the location of anomalies along with the anomaly type information, which results in the generated anomalies only distributed in specific locations. Therefore, we propose spatial anomaly embedding $e$, consisting of an anomaly embedding $e_{a}$ (for appearance) and a spatial embedding $e_{s}$ (for location), which disentangles the spatial information from anomaly appearance. To further validate this theory, we directly employ text inversion to train an anomaly embedding and use the trained embedding to generate anomalies with a given mask by blended latent diffusion (the same as our generation process). The generated results are shown in Fig. 6. It can be seen that the generated result by Textual Inversion tends to generate anomalies at the location the same as the training sample, which limits its application in anomaly generation where anomalies can be located at arbitrary positions.
D.2 Ablation on the rate of anomalies
We conduct additional experiments with the rate of anomalies as 10%, 20%, 30% (Ours), 40%, and 50% and test the performance on anomaly localization measured by AUROC and AP. The performance of anomaly localization is shown in Tab. 7. It can be seen that AP decreases quickly when the anomaly rate falls below 30%. This is attributed to the limited availability of training data for most categories, often comprising only 1-2 instances, making it challenging for the model to capture the anomaly information. Conversely, when the rate exceeds 30%, the model performance is similar. This indicates that our model can effectively learn sufficient anomaly information without a heavy reliance on an abundance of training samples.
Anomaly Rate 10% 20% 30% (Ours) 40% 50%
AUROC $\uparrow$ 96.2 98.1 99.1 99.0 98.7
AP $\uparrow$ 64.2 75.5 81.4 81.1 80.0
Table 7: ablation study on the rate of anomalies.
D.3 Ablation on the hyperparameters
We conduct ablation studies on the length of the anomaly embedding $l_{a}$ and spatial embedding $l_{s}$ in Tab. 8. Specifically, we train models with different $l_{a}$ and $l_{s}$ and then employ their generated data to train a UNet to localize the anomalies. It can be seen that when increasing $l_{s}$ and $l_{a}$, the total parameter number of the model rises, but the final performance in the downstream anomaly localization task is similar, which demonstrates that our model is not snesitive to the hyperparameters.
Model AUROC $\uparrow$ AP $\uparrow$ $F_{1}$-max $\uparrow$ PRO $\uparrow$
$L_{s}=4,l_{a}=8$ (Ours) 99.1 81.4 76.3 94.0
$L_{s}=4,l_{a}=16$ 98.8 80.6 75.1 93.2
$L_{s}=8,l_{a}=8$ 99.0 80.9 75.8 93.5
$L_{s}=8,l_{a}=16$ 99.1 81.2 75.9 93.8
Table 8: Ablation study on the anomaly embedding $l_{a}$ and spatial embedding $l_{s}$
D.4 Ablation on SAE and AAR
We conduct more ablation studies on the effectiveness of our spatial anomaly embedding (SAE) and adaptive attention re-weighting mechanism (AAR) by adding SAE and AAR separately. We compare our model with 3 models: 1) w/o AAR&SAE, 2) AAR only, and 3) SAE only in generating glue anomalies to the leather in the Fig. 7. It shows that the model without AAR&SAE cannot generate authentic anomalies or fill anomaly mask. While adding SAE improves anomaly authenticity, it doesn’t fill the mask. Moreover, incorporating AAR fills the mask but sacrifices authenticity. In contrast, our model (SAE + AAR) effectively generates authentic anomalies filling the mask.
Figure 7: Ablation study on SAE and AAR.
@@ Appendix G More Quantitative experiments
More comparison with the anomaly generation models. In this section, we provide supplementary experiments to complement those presented in the main paper. Specifically, in addition to the methods covered in the main paper, we include Crop&Paste (Lin et al. 2021) for comparison and we additionally introduce the Per Region Overlap (PRO) metric to provide a more comprehensive evaluation on anomaly localization. The experiment settings are the same as that in the main paper, where we train an U-net on the generated anomaly data. The pixel-level anomaly localization results are shown in Tab. 10 and the image-level anomaly detection results are shown in Tab. 11. The quantitative results demonstrate that our model outperforms all the other anomaly generation methods in terms of both anomaly localization and detection, indicating our good anomaly generation quality and diversity.
More comparison with the anomaly localization models. In this section, we further compare the anomaly detection methods with $F_{1}$-max score on anomaly localization. The results are shown in Table 12. It can be seen that our model achieves the best performance in anomaly localization with $F_{1}$-max.
@@
$F_{1}$-max KDAD CFLOW DREAM SSPCAB CFA RD4AD PatchCore DevNet DRA Ours
bottle 50.9 9.5 83.0 80.3 75.9 82.1 78.6 64.6 53.5 87.3
cable 18.3 10.5 58.5 51.2 76.3 65.2 68.5 54.9 55.3 83.5
capsule 15.1 6.9 48.9 49.5 57.0 60.4 56.7 38.7 47.7 60.8
carpet 54.2 3.5 60.0 47.1 48.3 67.8 67.9 52.3 42.3 74.6
grid 10.9 3.2 56.3 58.4 32.2 59.9 49.1 42.9 50.1 54.6
hazelnut 37.5 3.9 80.6 88.9 61.4 70.0 68.1 22.4 47.2 90.6
leather 30.4 3.9 63.2 58.1 53.8 67.2 54.7 32.1 19.8 71.0
metal nut 34.2 30.7 84.4 87.8 87.1 77.0 86.0 65.2 64.6 94.0
pill 29.9 17.6 62.6 46.5 79.5 63.7 73.5 22.8 45.5 90.8
screw 8.3 0.9 66.9 63.8 37.8 58.7 47.2 14.8 0.7 50.9
tile 27.8 26.6 90.8 88.5 77.8 71.8 69.4 69.9 61.4 86.2
toothbursh 25.1 4.7 47.5 37.2 62.1 58.7 63.8 35.1 22.6 73.4
transistor 26.7 19.9 55.2 34.8 76.3 59.8 64.2 28.9 33.2 85.7
wood 25.2 10.0 75.1 68.7 48.6 61.3 60.3 51.9 49.9 74.5
zipper 26.8 4.5 68.2 73.7 65.8 69.4 70.0 45.6 56.9 79.2
average 28.1 10.4 66.7 62.3 61.4 66.2 65.2 42.8 43.4 76.4
Table 12: Comparison on anomaly localization with $\mathbf{F_{1}}$-max.